# A possible connection between quantum mechanics and spacetime

### v3

@@ 2.2 The general system
In the last subsection, we show that for a free particle, the wave-particle duality can be interpreted as the uncertainty of the spacetime structure for the microscopic particle. It is well known that the non commutative relationship is one of the most fundamental equations in quantum mechanics. Both the wave-particle duality and the non commutative relationship hold true regardless of whether the particle is involved in interactions. In this subsection, we will show that in all scenarios, the wave-particle duality can always be interpreted as the uncertainty of the spacetime structure for the particle. And the path integral for any quantum system can be transformed into an integration over various spacetime metrics.
@@
$S_{tot}=\int_{\sigma_{i}}^{\sigma_{f}}d\sigma\big\{\sum_{n=1}^{N}m_{n}\sqrt{g_{\alpha\beta}\frac{dX^{\alpha}_{n}}{d\sigma}\frac{dX^{\beta}_{n}}{d\sigma}}+V(X^{\mu}_{1},...,X^{\mu}_{N})\big\}.$ (11)
Here, $N$ represents the total number of the particles. $\sigma$ is an arbitrary parameter. $X^{\mu}_{n}$ and $m_{n}$ represent the coordinate and the rest mass of the $n$-th particle, respectively. And $V(X^{\mu}_{1},...,X^{\mu}_{N})$ represents the interactions between different particles. Taking $\delta S_{tot}/\delta X^{\mu}_{n}=0$, one can obtain the classical dynamical equation of the system under general spacetime $g_{\alpha\beta}$ as
@@
Specifying the spacetime metric and the initial state of the system, equation (12) has only one solution. This solution represents the trajectories of the particles. We denote this solution as $\{X^{\mu}_{n}(\sigma)=\mathcal{K}^{\mu}_{n}(g_{\alpha\beta})\ |\ n=1,2,...,N\}$. Similar to the last subsection, intuitively speaking, changing the spacetime metric can lead the trajectories $\{X^{\mu}_{n}\ |\ n=1,2,...,N\}$ to become any set of curves (physically allowed). Thus, it is reasonable to deduce that by providing any set of curves (composed of $N$ physically allowed trajectories), one can find a class of metric $g_{\alpha\beta}]$ in which this set of curves serves as the solution to equation (12).
@@
According to conventional quantum mechanics in Minkowski spacetime ($g_{\alpha\beta}=\eta_{\alpha\beta}$), the transition amplitude of the system defined by equation (11) from the initial state $|X^{\mu}_{1}(\sigma_{i}),...,X^{\mu}_{N}(\sigma_{i})\rangle$ to the final state $|X^{\mu}_{1}(\sigma_{f}),...,X^{\mu}_{N}(\sigma_{f})\rangle$ is
@@
$\begin{split}&\ \ \ \langle X^{\mu}_{N}(\sigma_{f}),...,X^{\mu}_{1}(\sigma_{f})|X^{\mu}_{1}(\sigma_{i}),...,X^{\mu}_{N}(\sigma_{i})\rangle\\ &=\mathscr{N}\int\prod_{n=1}^{N}DX^{\mu}_{n}\ \mathrm{exp}\Big\{i\int_{\sigma_{i}}^{\sigma_{f}}d\sigma\big\{\sum_{n=1}^{N}m_{n}\sqrt{\eta_{\alpha\beta}\frac{dX^{\alpha}_{n}}{d\sigma}\frac{dX^{\beta}_{n}}{d\sigma}}+V(X^{\mu}_{1},...,X^{\mu}_{N})\big\}\Big\}.\end{split}$ (13)
@@
The path integral $\int\prod_{n=1}^{N}DX^{\mu}_{n}$ represents that all possible sets of trajectories $\{X^{\mu}_{n}\ |\ n=1,2,...,N\}$ are summed over.
@@
We have argued that for any given set of trajectories $\{X^{\mu}_{n}\ |\ n=1,2,...,N\}$, it is always possible to embed them into a class of metric $g_{\alpha\beta}]$ where the trajectories satisfy equation (12). Therefore, the integration over different trajectories $\{X^{\mu}_{n}\ |\ n=1,2,...,N\}$ in equation (LABEL:eq:1.12) can be transformed into an integration over various classes of spacetime metric $g_{\alpha\beta}]$. Following the similar statements of the last subsection, the path integral (LABEL:eq:1.12) can be transformed into
@@
$\begin{split}&\ \ \ \langle X^{\mu}_{N}(\sigma_{f}),...,X^{\mu}_{1}(\sigma_{f})|X^{\mu}_{1}(\sigma_{i}),...,X^{\mu}_{N}(\sigma_{i})\rangle\\ &=\mathscr{N}\int Dg_{\mu\nu}\ \mathcal{J}_{N}\ \Delta(\mathcal{G}^{(m)}_{N})\prod_{m=1}^{M_{N}}\delta(\mathcal{G}^{(m)}_{N})\ \\ &\ \ \ \ \ \times\mathrm{exp}\Big\{i\int_{\sigma_{i}}^{\sigma_{f}}d\sigma\big\{\sum_{n=1}^{N}m_{n}\sqrt{g_{\alpha\beta}\frac{d\mathcal{K}^{\alpha}_{n}}{d\sigma}\frac{d\mathcal{K}^{\beta}_{n}}{d\sigma}}+V(\mathcal{K}^{\mu}_{1},...,\mathcal{K}^{\mu}_{N})\big\}\Big\}.\end{split}$ (14)
@@
Here, $\mathcal{J}_{N}$, $\Delta(\mathcal{G}^{(m)}_{N})$ and $\mathcal{G}^{(m)}_{N}$ represent the Jacobian determinant, the Faddeev-Popov determinant and the gauge conditions, respectively. $M_{N}$ represent the number of the gauge conditions. These quantities differ from those of a free particle. Once again, we refrain from explicitly determining their forms in this study.
@@
Equation (LABEL:eq:1.13) is equivalent to equation (LABEL:eq:1.12). Equation (LABEL:eq:1.13) indicates that in a general system, integrating over different trajectories can be transformed into integrating over various spacetime metrics $g_{\alpha\beta}$. In equation (LABEL:eq:1.13), the spacetime metric is the only integral variable. Thus, the wave-particle duality can generally be interpreted as the uncertainty of the spacetime structure for microscopic particle. Therefore, it is reasonable to infer that the quantum characteristics of a general system can be ascribed to the uncertainty of the spacetime structure. Moreover, this perspective implies a potential connection between quantum characteristics and geometric quantities. It is worth noting that the Ryu-Takayanagi (RT) formula indicates that the entanglement entropy of a conformal field theory is equal to the area of a minimal surface [24, 34]. The RT formula thus serves as a specific example illustrating the relationship between quantum characteristics and geometric quantities.

### v4

@@ 2.2 The General System
In the last subsection, we show that for a free particle, wave–particle duality can be interpreted as the uncertainty of the spacetime structure for the microscopic particle. It is well known that the non-commutative relationship is one of the most fundamental equations in quantum mechanics. Both wave–particle duality and the non-commutative relationship hold true regardless of whether the particle is involved in interactions. In this subsection, we will show that in all scenarios, wave–particle duality can always be interpreted as the uncertainty of the spacetime structure for the particle. Furthermore, the path integral for any quantum system can be transformed into an integration over various spacetime metrics.
@@
$S_{tot}=\int_{\sigma_{i}}^{\sigma_{f}}d\sigma\big\{\sum_{n=1}^{N}m_{n}\sqrt{g_{\alpha\beta}\frac{dX^{\alpha}_{n}}{d\sigma}\frac{dX^{\beta}_{n}}{d\sigma}}+V(X^{\mu}_{1},\ldots,X^{\mu}_{N})\big\}.$ (10)
Here, $N$ represents the total number of the particles. $\sigma$ is an arbitrary parameter. $X^{\mu}_{n}$ and $m_{n}$ represent the coordinate and the rest mass of the $n$-th particle, respectively. Furthermore, $V(X^{\mu}_{1},\ldots,X^{\mu}_{N})$ represents the interactions between different particles. Taking $\delta S_{tot}/\delta X^{\mu}_{n}=0$, one can obtain the classical dynamical equation of the system under general spacetime $g_{\alpha\beta}$ as
@@
Specifying the spacetime metric and the initial state of the system, Equation (11) has only one solution. This solution represents the trajectories of the particles. We denote this solution as $\{X^{\mu}_{n}(\sigma)=\mathcal{K}^{\mu}_{n}(g_{\alpha\beta})\ |\ n=1,2,\ldots,N\}$. Similar to the last subsection, intuitively speaking, changing the spacetime metric can lead the trajectories $\{X^{\mu}_{n}\ |\ n=1,2,\ldots,N\}$ to become any set of curves (physically allowed). Thus, it is reasonable to deduce that by providing any set of curves (composed of $N$ physically allowed trajectories), one can find a class of metrics $g_{\alpha\beta}]$ in which this set of curves serves as the solution to Equation (11).
@@
According to conventional quantum mechanics in Minkowski spacetime ($g_{\alpha\beta}=\eta_{\alpha\beta}$), the transition amplitude of the system defined by Equation (10) from the initial state $|X^{\mu}_{1}(\sigma_{i}),\ldots,X^{\mu}_{N}(\sigma_{i})\rangle$ to the final state $|X^{\mu}_{1}(\sigma_{f}),\ldots,X^{\mu}_{N}(\sigma_{f})\rangle$ is
@@
$\begin{split}&\ \ \ \langle X^{\mu}_{N}(\sigma_{f}),\ldots,X^{\mu}_{1}(\sigma_{f})|X^{\mu}_{1}(\sigma_{i}),\ldots,X^{\mu}_{N}(\sigma_{i})\rangle\\ &=N\int\prod_{n=1}^{N}DX^{\mu}_{n}\ \mathrm{exp}\Big\{i\int_{\sigma_{i}}^{\sigma_{f}}d\sigma\big\{\sum_{n=1}^{N}m_{n}\sqrt{\eta_{\alpha\beta}\frac{dX^{\alpha}_{n}}{d\sigma}\frac{dX^{\beta}_{n}}{d\sigma}}+V(X^{\mu}_{1},\ldots,X^{\mu}_{N})\big\}\Big\}.\end{split}$ (12)
@@
The path integral $\int\prod_{n=1}^{N}DX^{\mu}_{n}$ represents that all possible sets of trajectories $\{X^{\mu}_{n}\ |\ n=1,2,\ldots,N\}$ are summed over.
@@
We have argued that for any given set of trajectories $\{X^{\mu}_{n}\ |\ n=1,2,\ldots,N\}$, it is always possible to embed them into a class of metrics $g_{\alpha\beta}]$ where the trajectories satisfy Equation (11). Therefore, the integration over different trajectories $\{X^{\mu}_{n}\ |\ n=1,2,\ldots,N\}$ in Equation (LABEL:eq:1.12) can be transformed into an integration over various classes of spacetime metrics $g_{\alpha\beta}]$. Following the similar statements of the last subsection, the path integral (LABEL:eq:1.12) can be transformed into
@@
$\begin{split}&\ \ \ \langle X^{\mu}_{N}(\sigma_{f}),\ldots,X^{\mu}_{1}(\sigma_{f})|X^{\mu}_{1}(\sigma_{i}),\ldots,X^{\mu}_{N}(\sigma_{i})\rangle\\ &=N\int Dg_{\mu\nu}\ \mathcal{J}_{N}\ \Delta(\mathcal{G}^{(m)}_{N})\prod_{m=1}^{M_{N}}\delta(\mathcal{G}^{(m)}_{N})\ \\ &\ \ \ \ \ \times\mathrm{exp}\Big\{i\int_{\sigma_{i}}^{\sigma_{f}}d\sigma\big\{\sum_{n=1}^{N}m_{n}\sqrt{g_{\alpha\beta}\frac{d\mathcal{K}^{\alpha}_{n}}{d\sigma}\frac{d\mathcal{K}^{\beta}_{n}}{d\sigma}}+V(\mathcal{K}^{\mu}_{1},\ldots,\mathcal{K}^{\mu}_{N})\big\}\Big\}.\end{split}$ (13)
@@
Here, $\mathcal{J}_{N}$, $\Delta(\mathcal{G}^{(m)}_{N})$ and $\mathcal{G}^{(m)}_{N}$ represent the Jacobian determinant, the Faddeev–Popov determinant and the gauge conditions, respectively. $M_{N}$ represents the number of gauge conditions. These quantities differ from those of a free particle. Once again, we refrain from explicitly determining their forms in this study.
@@
Equation (LABEL:eq:1.13) is equivalent to Equation (LABEL:eq:1.12). Equation (LABEL:eq:1.13) indicates that in a general system, integrating over different trajectories can be transformed into integrating over various spacetime metrics $g_{\alpha\beta}$. In Equation (LABEL:eq:1.13), the spacetime metric is the only integral variable. Thus, wave–particle duality can generally be interpreted as the uncertainty of the spacetime structure for microscopic particles. Therefore, it is reasonable to infer that the quantum characteristics of a general system can be ascribed to the uncertainty of the spacetime structure. Moreover, this perspective implies a potential connection between quantum characteristics and geometric quantities.